OpenStudy (teller):

Mind telling me If I got this right? An amount of $49,000 is borrowed for 13years at 7.5% interest, compounded annually. If the loan is paid in full at the end of that period, how much must be paid back?Use the calculator provided and round your answer to the nearest dollar. f=49000(1+0.075)^13 =1254602.40378726 or f=49000(1-0.075)^13= 177843.74932331

OpenStudy (solomonzelman):

\(\LARGE\color{black}{ \displaystyle \color{blue}{\rm A}=\color{royalblue}{\rm P}\left(1+\frac{ \color{red}{\rm r}}{\color{green}{\rm n}} \right)^{\huge \color{green}{\rm n} \times\color{magenta}{\rm t}} }\) \(\Large \color{black}{ \displaystyle \color{blue}{\rm A}}\) - is the final amount after the investment. \(\Large \color{black}{ \displaystyle \color{royalblue}{\rm P} }\) - is the investment from which you started. \(\Large \color{black}{ \displaystyle \color{red}{\rm r} }\) - is the percent rate (converted to decimal) \(\Large \color{black}{ \displaystyle \color{green}{\rm n} }\) - the number of times that interest is compounded each year. \(\Large \color{black}{ \displaystyle \color{magenta}{\rm t} }\) - the number of years for which you invest.

OpenStudy (solomonzelman):

Do you aslo see stuff crossed out?

OpenStudy (teller):

I already did that...

OpenStudy (solomonzelman):

tell me, do you see normal latex in olors, or crossed out codes?

OpenStudy (solomonzelman):

colors*

OpenStudy (teller):

i just see normal colors.

OpenStudy (solomonzelman):

oh, cool, so use that formula.

OpenStudy (teller):

I did use that formula just in a different format the one we were told to use..

OpenStudy (teller):

f=p(1+r)^1 or f=p(1-r)^t

OpenStudy (solomonzelman):

An amount of $49,000 is borrowed for 13years at 7.5% interest, compounded annually. If the loan is paid in full at the end of that period, how much must be paid back?Use the calculator provided and round your answer to the nearest dollar. t=13 r=0.075 n=1 P=49000 "A" is what you need to find.

OpenStudy (solomonzelman):

Inside should be a +

OpenStudy (teller):

yea and i already found them but i need to know which one is right.

OpenStudy (solomonzelman):

I have a wierd glitch, don't know why. but as long as you can see everything fine, it is not that bad

OpenStudy (solomonzelman):

P(1+r)^(t)

OpenStudy (solomonzelman):

that one

OpenStudy (teller):

Thank you so how do i round it ?

OpenStudy (solomonzelman):

did you find the value?

OpenStudy (teller):

1254602.40378726 is what i got i just need to know how to round it now.

OpenStudy (phi):

when you owe interest, it is added to the total cost. the final amount will be more than what you borrowed. that is how to remember (1+r) (and not (1-r) )

OpenStudy (solomonzelman):

well, the amount you will have to return is 125460.24

OpenStudy (teller):

It has to be rounded.

OpenStudy (solomonzelman):

well rounded to nearest cent.

OpenStudy (teller):

it said dollar...

OpenStudy (teller):

that's what i need help with.

OpenStudy (solomonzelman):

Well: 31, 32, 33, 34 would round up to 30 35, 36, 37, 38, 39 would round up to 40

OpenStudy (solomonzelman):

So from 5 to 9, rounds up. And from 1 to 4, rounds down

OpenStudy (teller):

1254602.3?

OpenStudy (solomonzelman):

you keep missplacing the point

OpenStudy (solomonzelman):

decimal is after the 0

OpenStudy (solomonzelman):

And you have 125460.\(\color{red}{2}\)4 and 60.\(\color{red}{2}\) would round to 61.0 or 60.0 ?

OpenStudy (teller):

so the answer is 125460.24?

OpenStudy (solomonzelman):

well, you want to round to nearest dollar, so then not.

OpenStudy (teller):

dude what the hell?You keep telling me 125460.24

OpenStudy (teller):

all your doing is confusing me even more.

OpenStudy (solomonzelman):

125460.24 is the answer with cents.

OpenStudy (solomonzelman):

They ask you to round that up to dollar

OpenStudy (teller):

HOW !?!?

OpenStudy (teller):

i've been asking you for the last 5 minutes how?

OpenStudy (teller):

not asking for cents it's asking for dollars.

OpenStudy (solomonzelman):

60.2 where would the 2 round it to? UP (and you have 61.0) or DOWN (and you have 60.0). ----------------------------------- Recall that: 31, 32, 33, 34 would round up to 30 35, 36, 37, 38, 39 would round up to 40

OpenStudy (solomonzelman):

"Recall that" and below is an example

OpenStudy (teller):

125461.24?

OpenStudy (solomonzelman):

125460.24 ---> 125460.00

OpenStudy (teller):

so the answer is 125460.00?

OpenStudy (solomonzelman):

yes